# Magnetorheological Characterization of Blood Analogues Seeded with Paramagnetic Particles

### v1

@@ 2.1 Experimental setup and procedure
Two blood analogues were employed, one Newtonian and one viscoelastic25. The Newtonian analogue, NBa, was an aqueous solution of 52 wtDimethyl sulfoxide (DMSO) and the viscoelastic analogue, VBa, was fabricated by adding 100 ppm of Xanthan Gum to the Newtonian one. The density of both solutions was approximately the same, $\rho\approx 1048$ kg/m3, at 25C. Regarding the VBa’s elastic characteristics, we used a Capillary Breakup Extensional Rheometer (CaBER) coupled with a high-speed camera (Photron FASTCAM Mini UX100) with a set of lenses (Optem Zoom 70 X), to obtain images of the sample breakup under one-dimensional extensional flow, using the slow retraction method26. The images corresponding to the elasto-capillary regime were treated to obtain the evolution of the filament’s diameter with time and, through an exponential fit27, the fluid’s longest relaxation time was obtained: $\lambda_{\mathrm{VBa}}=0.350\pm 0.044$ ms, which is correspondent to a subtle elasticity (data and setup not shown here).
@@
Two magnetic particles were employed, the M270 and MyOne Carboxylic Acid Dynabeads (Thermo Fisher Scientific). The M270, the larger of the two, had a diameter of $d\mathrm{p}=2.8$ µm while the MyOne only 1 µm. Grob et al. present magnetisation curves for both particles for magnetic field densities $-75\lesssim B\lesssim 75$ mT, and saturation magnetization values for $B\gtrsim 500$ mT: 6.4 and 9.4 Am2/kg for the M270 and MyOne, respectively. The particles were seeded in either fluid, being re-dispersed prior to all measurements to counteract sedimentation and break up any previously established structure. The M270 and MyOne particles had densities of 1600 and 1800 kg/m3, respectively.
@@
The sample was carefully loaded onto the geometry’s inferior plate with a VWR® standard line precision pipette. The steady shear measurements ranged shear rates $1\leq\dot{\gamma}\leq 10000$ s-1, with a constant time interval of 20 s between shear rate shift and torque measurement. The oscillatory shear response was evaluated by performing amplitude sweeps between $0.1\leq\gamma_{0}\leq 100$ at frequencies between $0.1\leq\omega\leq 2$ rad/s.
@@
The magnetic particles are a precious resource, and the number of measurements was minimised. For steady shear, a minimum of two repetitions was deemed acceptable, having that both measurements agreed reasonably well; if significant data discrepancies were found additional measurements were conducted up to a maximum of five.
@@
A compact experimental campaign allowed us to verify the effects of the magnetic field density, particle concentration, particle type and continuous phase (NBa or VBa). We selected benchmark measurements from which relevant characteristics could be independently altered. These were samples of 15 wtof the MyOne particles under the influence of a magnetic field density of 250 mT. We opted to test the continuous phase more in-depth, i.e., the blood analogues; thus, the variation of the remaining properties was tested both on NBa and VBa samples. To test for the influence of the field density, it was reduced from the benchmark 250 mT to 50 mT. Similarly, testing the particle concentration, it was reduced from 15 wtto 5 wt(volume fractions of approximately: 3.4 and 10.6% for the M270, and 3.0 and 9.7% for the MyOne particles). Lastly, the seeded particles were the M270 instead of the MyOne. This results, for either analogue as the continuous phase, on the following measurements, with their respective characteristic of interest and nomenclature, in parenthesis, to be used hereafter:
@@
Unseeded 0 wt- 0 mT Analogue rheology
15_MyOne_250 15 wtMyOne 250 mT Benchmark
5_MyOne_250 5 wtMyOne 250 mT Particle concentration
15_MyOne_50 15 wtMyOne 50 mT Magnetic field density
15_M270_250 15 wtM270 250 mT Particle type
@@ 3 Results and discussion
Previous steady shear measurements
Initially, we employed a smooth Cone-and-Plate geometry for our measurements (CP20 MRD). However, this geometry showed clear signs of apparent slip, a well-known phenomenon associated with suspension rheology where a particle-depletion layer is formed near the geometry’s smooth walls and in which a large part of the imposed deformation is applied solely on the low-viscosity continuous phase29, 30. The serrated PP20 MRD P2 geometry seemingly rectified the apparent slip, but our initial testing also showed that this geometry is affected by a gap-error of approximately $\varepsilon=123$ µm. As such, all data gathered with the PP20 MRD P2 is shown corrected for this gap-error according to the formulation of Kramer et al. Moreover, we also gathered that low and high shear limitations constrict our experimental window from this initial testing. At low shear, the rheometer’s minimum torque limit32 is outside our experimental range. Still, surface tension forces acting along the sample’s contact line33 are responsible for uncertainties that can be estimated by a minimum torque 70 times larger than the instrument’s specification ($70\times\mathrm{T}_{\mathrm{min}}$). At high shear, sample inertia may lead to an apparent shear-thickening32 which is reasonably predicted by the secondary flow limit34. For conciseness, this preliminary data is not shown here, but a brief discussion is presented as supplementary material (S1).
@@ 3.1 Steady shear
Figure 1 shows the obtained viscosity curves for the different working fluids in Table 1. The columns of Figure 1 separate the samples with either analogue as the continuous phase, while the three rows compare the effects of altering the magnetic field density, the particle concentration and the particle type with the benchmark measurements and the unseeded analogues. Figure 1 also depicts the secondary flow34 and adjusted low-torque32 limits, which are given by Equations 1 and 2, respectively. Because the secondary flow limit is dependent on the gap height and the gap-error is significant, this limit was accordingly corrected ($h=h_{\mathrm{c}}+\varepsilon$).
$\eta>\frac{\rho\dot{\gamma}h^{3}}{4R}\,,$ (1)
$\eta>\frac{2\,(70\times\mathrm{T}_{\mathrm{min}})}{\pi R^{3}\dot{\gamma}}\,.$ (2)
Figure 1: Viscosity curves obtained with the NBa (left column) and VBa (right column), seeded with MyOne particles at 15 wtunder the influence of 50 and 250 mT (top row), with MyOne particles at 5 and 15 wtunder 250 mT (middle row), and with MyOne and M270 particles at 15 wtunder 250 mT (bottom row). Adjusted low-torque and secondary flow limits are also shown in each graph.
Figure 2: Flow curves obtained with the NBa (left column) and VBa (right column), seeded with MyOne particles at 15 wtunder the influence of 50 and 250 mT (top row), with MyOne particles at 5 and 15 wtunder 250 mT (middle row), and with MyOne and M270 particles at 15 wtunder 250 mT (bottom row). Adjusted low-torque and secondary flow limits are also shown in each graph..
@@
Reducing the magnetic field density from 250 to 50 mT did not result in a significant rheological alteration, but a slight overall viscosity reduction is noted. From the magnetization curves presented by Grob et al. we can gather that the particle magnetization is more significantly dependent on the field density for weaker fields, which could explain the subtle rheological alterations between 250 and 50 mT. On the other hand, reducing the particle concentration from 15 to 5 wtdid lead to a considerable reduction in low-shear viscosity, with the 5 wtsamples behaving almost similar to the unseeded VBa. Regarding the particle type, the M270 are significantly larger (2.8 times larger than the MyOne), but the role of particle size is not yet clear35. However, the M270 have a lesser magnetic susceptibility28 which is possibly the responsible parameter for the weaker magnetorheological response compared to the viscosity curve provided by the MyOne particles for the same concentration and magnetic field.
MR fluids are typically expected to have a yield stress. In Figure 2 are plotted the flow curves (shear stress, $\tau$, against shear rate), but for none of the tested samples does the stress converge at low shear. Moreover, a low-shear viscosity plateau is also not observable. As such, the present data does not allow either for an estimation of yield stresses nor zero-shear viscosities.
@@ 3.2.1 Unseeded analogues
Figure 3 shows the Pipkin diagrams36 of Lissajous curves and the first harmonic viscoelastic moduliiiiModuli data was deleted if specific error messages returned from the rheometer software regarding either the viscous or elastic torque, respectively. for the unseeded samples. The elastic (stress vs strain, in blue) and viscous (stress vs strain rate, in red) Lissajous curves show the raw stress response and are max-normalised (adimensional). The maximum measured stress is presented above the plots and the decoupled elastic and viscous stresses20 are also lightly plotted.
@@
Figure 3: Oscillatory shear data gathered with the unseeded NBa (left) and VBa (right), with the PP20 MRD P2. (Top) Pipkin diagrams depicting the Lissajous curves. The blue and red curves represent, respectively, the elastic (stress vs strain) and viscous (stress vs strain rate) Lissajous curves. The elastic and viscous stresses are also lightly plotted in the respective colours and the maximum measured stress is presented above each plot. (Bottom) First harmonic loss ($G^{{}^{\prime\prime}}_{1}$, in circular markers) and storage ($G^{{}^{\prime}}_{1}$, in triangular markers) moduli and experimental limits associated with low-torque issues and instrument and sample inertia (corrected for the geometry’s gap-error).
Focusing on the Lissajous curves (top of Figure 3), the data is clearly ridden with experimental error. Having that the NBa is Newtonian, the expected response would be a perfectly circular elastic curve and a positive-diagonal of the viscous response. At low frequency and moderate strain ($\omega=0.1$ rad/s, $1\leq\gamma_{0}\leq 10$) the NBa seems to present a viscoelastic response that could be due to additional torque from surface tension forces while at large frequency and amplitude ($1\leq\omega\leq 2$ rad/s, $\gamma_{0}=100$) the stress is in opposite phase with the strain rate, which could be a symptom of severe inertial effects. With the VBa, at low frequency and large strain ($0.1\leq\omega\leq 0.5$ rad/s, $\gamma_{0}=100$) there is a semblance of the expected subtle viscoelasticity, but the remaining data is so deeply affected by experimental error that we cannot ascertain if this is representative of the fluid’s actual rheology.
As we have discussed, the first harmonic moduli provide limited information, but they are still useful to plot against known experimental limits, allowing to evaluate the measurement window. The low-torque limit, for a Parallel-Plate geometry, is given by32:
@@
where $G$ is either the loss ($G^{\prime\prime}$) or storage modulus ($G^{\prime}$), $\mathrm{T}_{\mathrm{min}}$ is the instrument’s minimum torque in oscillatory shear and $R$ is the geometry radius. At high frequencies, inertial limitations become important. On the one hand, the instrument inertia limit is given by32:
@@
where $I$ is the geometry’s moment of inertia and $h$ is the gap between plates. On the other hand, the sample inertia limit can be written as32:
@@
where $|G^{*}|=\sqrt{G^{\prime 2}+G^{\prime\prime 2}}$ is the magnitude of the complex modulus. Similar to the secondary flow limit for steady shear, we correct the inertia limits for the gap-error of the PP20 MRD P2. These experimental limits are also shown in Figure 3 (bottom). As can be seen, we are working with a quite narrow experimental window. On the one hand, we approach the low-torque limit at small amplitudes, which is not corrected by any safety coefficient to account for surface tension torque (contrarily to the steady-shear case). At high frequencies, inertial issues, also enlarged by the gap-error, become problematic with the instrument inertia limit always practically overrun. Of course, with the poor quality of these first measurements, little information can be gathered regarding which limitation is responsible for which result.
@@
What we can conclude, however, is that the PP20 MRD P2 is unsuitable for characterising the unseeded analogues’ rheological properties under oscillatory shear, at least within the tested experimental range. Looking at the expressions for the experimental limitations, the most helpful way to widen the experimental window is to increase the geometry diameter. Thus, to get a glimpse of the actual fluid responses, we performed the same oscillatory shear measurements with a standard, smooth PP50 in an additional inferior plate (the magnetorheological cell has a geometrical constraint of approximately 30 mm on standard geometries and 20 mm on the MRD geometries). The obtained Pipkin diagrams and moduli plots are shown in Figure 4.
(PP50) Unseeded
Figure 4: Oscillatory shear data gathered with the unseeded (left) NBa and (right) VBa, with a standard, smooth PP50. (Top) Pipkin diagrams depicting the Lissajous curves. The blue and red curves represent, respectively, the elastic (stress vs strain) and viscous (stress vs strain rate) Lissajous curves. The elastic and viscous stresses are also lightly plotted in the respective colours and the maximum measured stress is presented above each plot. (Bottom) First harmonic loss ($G^{{}^{\prime\prime}}_{1}$, in circular markers) and storage ($G^{{}^{\prime}}_{1}$, in triangular markers) moduli and experimental limits associated with low-torque issues and instrument and sample inertia.
Analysing the Pipkin diagrams, we can now see the expected fluid responses at high deformation ($\gamma_{0}=100$). The NBa returned practically perfectly circular elastic Lissajous curves and the viscous curve was the expected positive-diagonal. Moreover, the elastic and viscous stress plots also show practically straight lines, which all indicate the purely viscous linear response we expect of a Newtonian fluid. On the other hand, at this imposed strain we can see the VBa presenting a slight bending of the viscous Lissajous, an indicator of additional harmonics emerging in the NLVE. Reducing the deformation to $\gamma_{0}=10$, the NBa shows some apparent viscoelastic behaviour at small frequencies, which, looking at the moduli plots, could be due to surface-tension torque. As such, this could also be the cause for the VBa’s viscoelastic response (at $\gamma_{0}=10$ and $\omega<1$ rad/s) despite its moduli being slightly larger. Nevertheless, at this amplitude, increasing the frequency still somewhat returns the NBa’s purely viscous behaviour (despite some oscillations possibly due to inertial issues), while the VBa presents a slight viscoelasticity which agrees with its very subtle elastic nature ($\lambda_{\mathrm{CaBER}}=0.35$ ms), and there is observable bending of the elastic stress which indicates we are still obtaining a non-linear response. Considering lower strains, $\gamma_{0}<10$, because the NBa data is no longer representative of a purely viscous behaviour we are wary of gathering conclusions of its viscoelastic counterpart. As such, we assume the data for $\gamma_{0}<10$ is no longer representative of either fluid’s rheology.
As discussed, there is evidence of surface tension torque at low frequency up to middling amplitudes (a non-negligible apparent elasticity of the NBa). Looking at the low-torque limit, we can presume, thus, that surface tension has indeed shifted the limit considerably and the data gathered at low frequency and amplitude should be viewed with scepticism. Considering inertial effects, it is difficult to lay the fault to the instrument or sample inertia, but errors are still dominating at high frequencies and small amplitudes. Despite these issues, the increase in diameter, possibly aided by the elimination of the gap-error (the estimated PP50 gap-error is practically insignificant37, $\varepsilon\approx 3$ µm), successfully widened the experimental window, allowing us to visualise the analogues’ response to oscillatory shear. In the current state, this is useful for standard measurements with samples of subtle rheological properties.

### v2

@@ 2.2 Working fluids
Two blood analogues were employed, one Newtonian and one viscoelasticcampo2013. The Newtonian analogue, NBa, was an aqueous solution of 52 wt% Dimethyl sulfoxide (DMSO) and the viscoelastic analogue, VBa, was fabricated by adding 100 ppm of Xanthan Gum to the Newtonian one. The density of both solutions was approximately the same, $\rho\approx 1050$ kg/m3, at 22.7C. Regarding the VBa’s elastic characteristics, we used a Capillary Breakup Extensional Rheometer (CaBER) coupled with a high-speed camera (Photron FASTCAM Mini UX100) with a set of lenses (Optem Zoom 70 X), to obtain images of the sample breakup under one-dimensional extensional flow, using the slow retraction methodcampo2010slow. The images corresponding to the elasto-capillary regime were treated to obtain the evolution of the filament’s diameter with time and, through an exponential fitmckinley2005, the fluid’s longest relaxation time was obtained: $\lambda_{\mathrm{VBa}}=0.350\pm 0.044$ ms, which is correspondent to a subtle elasticity (data and setup not shown here for brevity).
2.3 Magnetic particles
@@
Two magnetic particles were employed, the M270 and MyOne Carboxylic Acid Dynabeads (Thermo Fisher Scientific). The M270, the larger of the two, had a diameter of $d_{\mathrm{p}}=2.8$ µm while the MyOne only 1 µm. grob2018 present magnetisation curves for both particles (saturation magnetization values of $M_{\mathrm{s}}=9.4$ and 6.7 Am2/kg for the MyOne and M270, respectively, for $B\gtrsim 0.5$ T) The particles were seeded in either fluid, being re-dispersed prior to all measurements to counteract sedimentation and break up any previously established structure. The M270 and MyOne particles had densities of 1600 and 1800 kg/m3, respectively.
2.4 Experimental procedure
@@
The sample was carefully loaded onto the geometry’s bottom plate with a VWR® standard line precision pipette. The steady shear measurements ranged shear rates $1\leq\dot{\gamma}\leq 10000$ s-1, with a constant time interval of 20 s between shear rate shift and torque measurement. The oscillatory shear response was evaluated by performing amplitude sweeps between $0.1\leq\gamma_{0}\leq 100$ at frequencies between $0.1\leq\omega\leq 2$ rad/s.
@@
The magnetic particles are a precious resource, and the number of measurements was minimised. For steady shear, a minimum of two repetitions was deemed acceptable, having that both measurements agreed reasonably well; if significant data discrepancies were found, additional measurements were conducted up to a maximum of five.
@@
A compact experimental campaign allowed us to investigate the effects of the magnetic field density, particle concentration, particle type and continuous phase (NBa or VBa). We selected benchmark measurements that allowed for the independent variation of relevant characteristics. The samples consisted of 15 wt% of MyOne particles subjected to a magnetic field density of 250 mT. We focused particularly on testing the continuous phase, specifically the blood analogues, while varying the other properties for both NBa and VBa samples. To test for the influence of the field density, it was reduced from the benchmark 250 mT to 50 mT. Similarly, testing the particle concentration, it was reduced from 15 wt% to 5 wt% (volume fractions of approximately: 10.6 and 3.4% for the M270, and 9.7 and 3.0% for the MyOne particles). Lastly, we tested the M270 particles instead of the MyOne particles. Table 1 summarizes the nomenclature and the description of the measurements in relation to the seeded particle concentration, type and applied magnetic field density for both blood analogues as the continuous phase.
@@
Unseeded 0 wt% - 0 mT Analogue rheology
15_MyOne_250 15 wt% MyOne 250 mT Benchmark
5_MyOne_250 5 wt% MyOne 250 mT Particle concentration
15_MyOne_50 15 wt% MyOne 50 mT Magnetic field density
15_M270_250 15 wt% M270 250 mT Particle type
@@ 3 Results and discussion
Initially, we employed a smooth Cone-and-Plate geometry for our measurements (CP20 MRD). However, this geometry showed clear signs of apparent slip, a well-known phenomenon associated with suspension rheology, where a particle-depletion layer forms near the geometry’s smooth walls and in which a large part of the imposed deformation is applied solely to the low-viscosity continuous phasebarnes1995, vicente2004. The serrated PP20 MRD P2 geometry seemingly rectified the apparent slip, but our initial testing also showed that this geometry is affected by a gap-error of approximately $\varepsilon=123$ µm. Such errors are expected of detailed geometries because the sample flows within the geometry grooves, similar to a porous mediumcarotenuto2013, nickerson2005. As such, all data gathered with the PP20 MRD P2 are corrected for this gap-error according to the formulation of kramer1987, similarly to our previous workrodrigues2025. Moreover, we also gathered that low and high shear limitations constrain our experimental window from this initial testing. At low shear, the rheometer’s minimum torque limitewoldt2015 is outside our experimental range. Still, surface tension forces acting along the sample’s contact linejohnston2013 are responsible for uncertainties that can be estimated by a minimum torque 70 times larger than the instrument’s specification ($70\times\mathrm{T}_{\mathrm{min}}$). At high shear, sample inertia may lead to an apparent shear-thickeningewoldt2015 which is reasonably predicted by the secondary flow limitturian1972. For conciseness, this preliminary data is not shown here, but a brief discussion is presented as supplementary material (Section S1).
@@ 3.1 Steady shear
Figure 1 shows the obtained viscosity curves for the different working fluids in Table 1. The columns of Figure 1 separate the samples with either analogue as the continuous phase, while the three rows compare the effects of altering the magnetic field density, the particle concentration and the particle type with the benchmark measurements and the unseeded analogues. Figure 1 also depicts the secondary flowturian1972 and adjusted low-torqueewoldt2015 limits, which are given by Equations 1 and 2, respectively. Because the secondary flow limit is dependent on the gap height and the gap-error is significant, this limit was accordingly corrected ($h=h_{\mathrm{c}}+\varepsilon$).
$\text{Secondary flow}:\;\eta>\frac{\rho\dot{\gamma}h^{3}}{4R}\,$, (1)
$\text{Adjusted minimum torque}:\;\eta>\frac{2\,(70\times\mathrm{T}_{\mathrm{min}})}{\pi R^{3}\dot{\gamma}}\,.$ (2)
Figure 1: Viscosity curves obtained with the NBa (left column) and VBa (right column), seeded with MyOne particles at 15 wt% under the influence of 50 and 250 mT (top row), with MyOne particles at 5 and 15 wt% under 250 mT (middle row), and with MyOne and M270 particles at 15 wt% under 250 mT (bottom row). Adjusted low-torque and secondary flow limits are also shown in each graph.
Figure 2: Flow curves obtained with the NBa (left column) and VBa (right column), seeded with MyOne particles at 15 wt% under the influence of 50 and 250 mT (top row), with MyOne particles at 5 and 15 wt% under 250 mT (middle row), and with MyOne and M270 particles at 15 wt% under 250 mT (bottom row). Adjusted low-torque and secondary flow limits are also shown in each graph.
@@
Reducing the magnetic field density from 250 to 50 mT did not result in a significant rheological alteration, but a slight overall viscosity reduction is noted. From the magnetization curves presented by grob2018 we can gather that the particle magnetization is more significantly dependent on the field density for weaker fields, which could explain the subtle rheological alterations between 250 and 50 mT. On the other hand, reducing the particle concentration from 15 to 5 wt% did lead to a considerable reduction in low-shear viscosity, with the 5 wt% samples behaving almost similarly to the unseeded VBa. Regarding the particle type, the M270 is significantly larger (2.8 times larger than the MyOne), but the role of particle size is not yet cleardeGans2000. However, the M270 have a lesser magnetic susceptibilitygrob2018, which is possibly the responsible parameter for the weaker magnetorheological response compared to the viscosity curve provided by the MyOne particles for the same concentration and magnetic field.
MR fluids are typically expected to have a yield stress. In Figure 2, the flow curves (shear stress, $\tau$, against shear rate) are plotted, but for none of the tested samples does the stress clearly converge at low shear. With the 5wt% MyOne samples and the unseeded VBa, there seems to be an apparent stress stabilization at low shear, but we believe this should be due to residual surface tension torque. This lack of yield stress evidence could mean that the particle concentration or the applied field strength are not sufficient to result in a yield behaviour, but a low-shear viscosity plateau is also not observable. We suspect that the apparent slip is responsible, even with the serrated geometry, and may be present on the smooth inferior plate. This is also corroborated by a small, characteristic bending of the flow curves. In any case, we can still estimate the yield stresses by fitting the Casson stress equation:
$\tau^{1/2}=\tau_{\mathrm{y}}^{1/2}+(\eta_{\infty}\dot{\gamma})^{1/2}\,$, (3)
where $\tau_{\mathrm{y}}$ and $\eta_{\infty}$ are the yield stress and the large-shear viscosity, both of which are dependent on the particle magnetisation and the volume fraction. The fits were performed through a least-squares method on trimmed data sets ($10\leq\dot{\gamma}\leq 7000$ s-1), avoiding possible errors from apparent slip at low-shear and secondary flows at high-shear. Table 2 gives the estimated yield stresses and low-shear viscosities, and the Casson model fitting curves are shown in the top row of Figure 3.
Table 2: Estimated yield stress and large-shear viscosity for each tested sample and magnetic field combination. Obtained through least-squares fit of the Casson model (Equation 3) to the trimmed stress data ($10\leq\dot{\gamma}\leq 7000$ s-1)
Continuous phase NBa VBa
[…]
15_MyOne_250 1.764 1.939 1.908 2.304
15_MyOne_50 0.880 2.028 0.922 2.429
5_MyOne_250 0.080 2.724 0.230 3.081
15_M270_250 0.529 2.304 0.666 2.914
The Casson model seems to reasonably fit the experimental data at moderate to large shear rates, but diverges at low-shear, $\dot{\gamma}\lesssim 10$ (even when fitting to the whole shear rate rangeiiiiiiThe Bingham model was also tested ($\tau=\tau_{\mathrm{y}}+\eta_{\infty}\dot{\gamma}$), but fit the data poorly for $\dot{\gamma}\lesssim 100$ s-1. Data shown in Supplementary Material, Section S2.), which was expected from the previously discussed lack of stress stabilisation. Nevertheless, the estimated yield stress values, despite quite low, seem to be representative of the magnetorheological strength, increasing with particle concentration and magnetic field density. Moreover, there is a general agreement between the tested continuous phases, with the VBa’s returning slightly larger yield stress and large-shear viscosity values than the NBa’s. The only significant discrepancy is found with the weakest magnetorheological samples, 5wt% of MyOne particles and 250 mT, which may be more prone to slip due to the lack of a percolating structure.
An alternative analysis can be achieved through the adaptation of the Casson plastic equation (Equation 3) considering the Mason number, $Mn$, and its critical value, $Mn^{*}$ (which determines the transition from the magnetostatic to the hydrodynamic regime):
$Mn=\frac{72\eta_{\mathrm{c}}\dot{\gamma}}{\mu_{0}\mu_{\mathrm{cr}}(\rho_{\mathrm{p}}M)^{2}}\,$, (4)
$Mn^{*}=\frac{72\eta_{\mathrm{c}}\tau_{\mathrm{y}}}{\mu_{0}\mu_{\mathrm{cr}}(\rho_{\mathrm{p}}M)^{2}\eta_{\infty}}\,.$ (5)
Here $\eta_{\mathrm{c}}$ is the continuous phase viscosity, $\mu_{0}$ is the vacuum permeability, $\mu_{\mathrm{cr}}$ is the continuous phase relative permeability ($\mu_{\mathrm{cr}}=1$), $\rho_{\mathrm{p}}$ is the particle density and $M$ is the particle magnetisation. A dimensionless form of the Casson equation can thus be written asmorillas2019, berli2012:
$\eta/\eta_{\infty}=1+(Mn/Mn^{*})^{-1}+2(Mn/Mn^{*})^{-1/2}\,.$ (6)
The bottom row of Figure 3 shows the obtained dimensionless viscosity curves along with the Casson model prediction given in Equation 6. The experimental data seems to reasonably collapse onto a single curve, as predicted by the Casson model. There are, however, significant deviations with the weakest magnetorheological tests (5 wt% of MyOne particles and 250 mT). At high $Mn$ the viscosity curves follow the expected behaviour but drop-off at $Mn/Mn^{*}\lesssim 1$. This effect is more clear with the VBa sample, whereas the NBa’s viscosity seems to re-collapse onto the Casson model for $Mn/Mn^{*}\lesssim 0.1$. In reality, this apparent viscosity increase ought to be related to low-shear errors, probably due to residual surface tension torque (see Figures 1 and 2). With the remaining samples, a similar viscosity drop-off is also observable at very low Mason numbers ($Mn/Mn^{*}\lesssim 0.03$), which could be due to apparent slip, as discussed before.
Figure 3: Flow curves with Casson model fits (top row) and dimensionless viscosity as a function of the reduced Mason number (bottom row), for each tested particle type, concentration and magnetic field density combination, for samples with the NBa (left) and VBa (right) as continuous phase (unseeded data not shown). Secondary flow limit is shown along with the flow curves (top row).
@@ 3.2.1 Unseeded analogues
Figure 4 shows the Pipkin diagramspipkin2012 of Lissajous curves and the first-harmonic viscoelastic moduliiiiiiiiiiModuli data was deleted if specific error messages returned from the rheometer software regarding either the viscous or elastic torque, respectively. for the unseeded samples. The elastic (stress vs strain, in blue) and viscous (stress vs strain rate, in red) Lissajous curves show the raw stress response and are max-normalised (adimensional). The maximum measured stress is presented above the plots and the decoupled elastic and viscous stressescho2005 are also lightly plotted.
@@
Figure 4: Oscillatory shear data gathered with the unseeded NBa (left) and VBa (right), with the PP20 MRD P2. (Top) Pipkin diagrams depicting the Lissajous curves. The blue and red curves represent, respectively, the elastic (stress vs strain) and viscous (stress vs strain rate) Lissajous curves. The elastic and viscous stresses are also lightly plotted in the respective colours and the maximum measured stress is presented above each plot. (Bottom) First-harmonic loss ($G^{\prime\prime}_{1}$, in circular markers) and storage ($G^{\prime}_{1}$, in triangular markers) moduli and experimental limits associated with low-torque issues and instrument and sample inertia (corrected for the geometry’s gap-error).
Focusing on the Lissajous curves (top of Figure 4), the data is clearly ridden with experimental error. Since the NBa is Newtonian, the expected response would be a perfectly circular elastic curve and a positive-diagonal of the viscous response. At low frequency and moderate strain ($\omega=0.1$ rad/s, $1\leq\gamma_{0}\leq 10$) the NBa seems to present a viscoelastic response that could be due to additional torque from surface tension forces while at large frequency and amplitude ($1\leq\omega\leq 2$ rad/s, $\gamma_{0}=100$) the stress is in opposite phase with the strain rate, which could be a symptom of severe inertial effects. With the VBa, at low frequency and large strain ($0.1\leq\omega\leq 0.5$ rad/s, $\gamma_{0}=100$) there is a semblance of the expected subtle viscoelasticity. Still, the remaining data is so deeply affected by experimental error that it is not possible to ascertain whether this represents the fluid’s actual rheology.
As we have discussed, the first-harmonic moduli provide limited information, but they are still useful to plot against known experimental limits, allowing to evaluate the measurement window. The low-torque limit, for a Parallel-Plate geometry, is given byewoldt2015:
@@
where $G$ is either the loss ($G^{\prime\prime}$) or storage modulus ($G^{\prime}$), $\mathrm{T}_{\mathrm{min}}$ is the instrument’s minimum torque in oscillatory shear and $R$ is the geometry radius. At high frequencies, inertial limitations become important. On the one hand, the instrument inertia limit is given byewoldt2015:
@@
where $I$ is the geometry’s moment of inertia and $h$ is the gap between plates. On the other hand, the sample inertia limit can be written asewoldt2015:
@@
where $|G^{*}|=\sqrt{G^{\prime 2}+G^{\prime\prime 2}}$ is the magnitude of the complex modulus. Similar to the secondary flow limit for steady shear, we correct the inertia limits for the gap-error of the PP20 MRD P2. These experimental limits are also shown in Figure 4 (bottom). As can be seen, the reliable experimental window is very small. On the one hand, we approach the low-torque limit at small amplitudes, which is not corrected by any safety coefficient to account for surface tension torque (contrarily to the steady-shear case). At high frequencies, inertial issues, also enlarged by the gap-error, become problematic with the instrument inertia limit always practically overrun. Therefore, little information can be gathered regarding which limitation is responsible for which result.
@@
What we can conclude, however, is that the PP20 MRD P2 is unsuitable for characterising the unseeded analogues’ rheological properties under oscillatory shear, at least within the tested experimental range. Looking at the expressions for the experimental limitations, the most helpful way to widen the experimental window is to increase the geometry diameter. We performed the same oscillatory shear measurements with a larger, smooth PP50 (Parallel-Plate with 50 mm of diameter) in an additional bottom plate (the magnetorheological cell has a geometrical constraint of approximately 30 mm on standard geometries and 20 mm on the MRD geometries) and found a significant improvement of the result quality. At large strain ($\gamma_{0}=100$) we encountered the expected purely-viscous NBa response, and evidence of non-linear viscoelasticity with the VBa. However, because our magnetorheological setup does not permit such measuring geometries, we present these results as supplementary material (Section S3).